OpenStudy (anonymous):

One endpoint of a segment is (12, -8). The midpoint is (3, 18). Find the coordinates of the other endpoint. Please help

hartnn (hartnn):

if u have 2 endpoints (x1,y1) and (x2,y2) with (m,n) as mid-point,then u have the relation m=(x1+x2)/2 n=(y1+y2)/2 from this,can u find (x2,y2)?

OpenStudy (anonymous):

well i orginally got (6, 44) for the last endpoint but wasnt sure

hartnn (hartnn):

thats incorrect,so what u get now?

OpenStudy (anonymous):

i dont understand how to use the equation

hartnn (hartnn):

ok, put m=3,x1=12 and find x2.....can u?

OpenStudy (anonymous):

i put in the numbers but i dont understand still

hartnn (hartnn):

let me do x2 for u,then u try for y2 3=(12+x2)/2 so,3*2=12+x2 so,6=12+x2 x2=6-12 x2=-6 got this?

OpenStudy (anonymous):

yeah

hartnn (hartnn):

so will u try for y2 and give me its value?

OpenStudy (anonymous):

i got -y2= -44

hartnn (hartnn):

18*2=-8+y2 36+8=y2 so y2=44 yup,u also got the same....

OpenStudy (anonymous):

so i got it right?

hartnn (hartnn):

yup,the point is (-6,44)

OpenStudy (anonymous):

okay... thank u for the help

hartnn (hartnn):

welcome :)